1. A solution of NaCl is neutral, with an expected PH of 7. If electrolysis was carried out on 500 mL of an NaCl solution with a current of 0.500 A, how many seconds would it take for the PH of the solution to rise to a value of 9? (PH is related to H+ concentration by PH = -log [H+]).
1
Expert's answer
2020-06-18T14:41:25-0400

Upon the electrolysis of the NaCl aqueous solution, the following processes take place:

anode: 2Cl- "\rightarrow" Cl2 + 2e-

cathode: 2H2O + 2e-"\rightarrow" H2 + 2OH-.

As one can see, hydroxide anions OH- are produced during the electrolysis. Therefore, the pH increases. At pH 9, the concentration of OH- anions is:

"c(OH^-) = 10^{-(14-pH)} = 10^{-5}" M.

This means that the number of the moles of OH- produced during the electrolysis is:

"n(OH^-) = cV = 10^{-5} (M)\u00b7500\u00b710^{-3}(L) = 5\u00b710^{-6}" mol.

The first Faraday's law states the relation between the number of the moles "n" of the substance produced during the electrolysis and the charge "Q" passed through the system:

"n = \frac{Q}{Fz}" ,

where "F" is the Faraday constant 96485 C/mol and "z" is the number of the electrons participating in the reaction divided by the stoichiometric coefficient of the substance produced. The hydroxide anion has a stoichiometric coefficient 2, so "z=1" . The charge passed through the solution is the product of the time and the current:

"Q = It" .

Therefore, the time needed to raise the pH to 9 is:

"t = \frac{nF}{I} = \frac{5·10^{-6}(mol)·96485(C/mol)}{0.500 (A)} = 0.96" s.

Answer: it would take 0.96 s for the pH of the solution to rise to a value of 9.